# FEM-based Neural Networks for Solving Incompressible Fluid Flows and Related Inverse Problems

### v1

@@ 2.3 Example 1: Stokes equations in 2D
[…]
Note that the homogeneous Neumann boundary condition on the outflow boundary automatically sets the mean outflow pressure to zero.
Figure 2: Triangular mesh around a NACA 0012 airfoil.
Homogeneous Neumann boundary conditions are known to be so-called natural conditions and thus included in (10) by construction. Nevertheless, after assembling the residual, it remains to enforce the inhomogeneous Dirichlet boundary conditions for the velocity. To do this, we exploit that the chosen basis functions are Lagrangian, i.e., each basis function $v_{h}^{i}$ has an associated Lagrangian node $x_{i}\in\Omega$ such that
$v_{h}^{i}(x_{j})=\delta_{ij}\quad\forall i,j.$ (13)
The Dirichlet boundary conditions are now enforced on all such $x_{j}\in\Omega^{+}\cup\Omega^{A}$ by adjusting the right hand side vector $\mathbf{f}$ and the stiffness matrix in (10) accordingly: First, we multiply the columns of the stiffness matrix corresponding to the boundary nodes by the specified boundary values and then subtract the result from $\mathbf{f}$. Next, we set the corresponding values of $\mathbf{f}$ to the boundary values. In the stiffness matrix we modify the rows and columns corresponding to the boundary nodes so that the diagonal value is unity and the off-diagonal entries are set to zero. Note that this is done before preconditioning, which is possible because the modifications do not affect the symmetry properties of the system (10).
2.4 Example 2: stationary Navier–Stokes equations in 2D
@@
i.e. compared to the Stokes equations (7)-(8) in particular the non-linear convection term $u(\lambda)\cdot\nabla u(\lambda)$ has been added. Moreover, we have introduced the so-called kinematic viscosity $\eta>0$ as additional parameter. Discretization leads to
@@
with $b$ as before, $a$ as before but multiplied by $\eta$, and the trilinear form $c(z_{h};u_{h},v_{h})=\int_{\Omega}(z_{h}\cdot\nabla u_{h})\cdot v_{h}$ originating from the nonlinear term. We again use Taylor-Hood elements for the discretization in the non-linear case, as these are already necessary for a stable solution in the linear case. To formulate the corresponding nonlinear algebraic equation, we utilize Lagrangian basis functions as in (10) and obtain the discrete residual
@@
i.e., $\mathbf{C}(\mathbf{u}_{\text{NN}})$ can be calculated by a contraction of the tensor $\mathbf{u}_{\text{NN}}$ with a 3D-tensor $\mathbf{\tilde{C}}$ that is independent of $\mathbf{u}$ and thus can be precomputed and reused. Since $\mathbf{\tilde{C}}$ is sparse, this can be done in a memory-efficient manner.
@@
$\mathbf{\hat{u}}_{i}:=\left\{\begin{array}[]{ll}\mathbf{u}_{i}&\text{if }x_{i}\notin\Omega^{+}\cup\Omega^{A},\\ u_{\textnormal{BC}}(x_{i})&\text{if }x_{i}\in\Omega^{+}\cup\Omega^{A},\end{array}\right.$ (19)
@@ 3 Numerical results
The respective neural networks and an FEM module as described in Sects. 2, 2.3, and 2.4 have been implemented in PyTorch [38]. Unless stated otherwise, the neural networks are plain vanilla feed forward networks consisting of five hidden layers with 50 neurons each and an output layer whose size is equal to the number of degrees of freedom of the discretization. Except for the last, linear, layer, we use SELU [39] as the activation function. To facilitate efficient training, we set the network weights with Xavier’s normal initialization [40] scaled by $3/4$. For the training, we use PyTorch’s implementation [^3] https://pytorch.org/docs/stable/generated/torch.optim.LBFGS.html [Accessed August 21, 2024] of L-BFGS. To make the results more comparable with “classical” implementations, we choose max_iter=1 to ensure that only one L-BFGS step is performed per iteration. Furthermore, we choose a line search within L-BFGS that seeks a solution that fulfills the strong Wolfe condition.
@@ 3.1 The non-parametric case
Both the preconditioned and the non-preconditioned NNs are trained and evaluated for a single angle of attack $\lambda=1^{\circ}$ and 40000 L-BFGS iterations, for each: the Stokes problem and the Navier–Stokes equation. In the latter case, the solutions are computed with viscosities $\eta=1$ and $\eta=0.1$. Figure 3 shows that in all these cases preconditioning significantly speeds up the training in terms of a drastically reduced number of iterations. For the preconditioned cases we can even observe a stagnation of the residual at levels that are probably related to numerical accuracy; this is probably related to a similar reason as the effect described in the paragraph below (12) in Sect. 2.3. It should be noted that training is also much faster in terms of runtime because the very significant reduction in the number of iterations more than offsets the minimal increase in the runtime of a single iteration that is due to the few additional operations related to preconditioning.
[…]
@@ 3.2 The parametric case
Figure 5 shows the loss values during training for the different configurations under consideration. Even in this case, preconditioning is able to accelerate the decay of the loss by more than two orders of magnitude for both the Stokes and the Navier–Stokes equations. For the Stokes problem, we also performed the experiments on a finer mesh, which is obtained from the original one by a single uniform refinement and results in 28085 degrees of freedom. The behavior of the losses on the original and the finer mesh can be taken as an indication of mesh independence, which is consistent with our expectations from Sect. A and our discussion in the previous section.
@@
Figure 5: Loss values per L-BFGS iteration when training FEM-based NNs for multiple angles of attack #$\#\lambda_{\textnormal{train}}$ with preconditioning (pre) and without preconditioning (nopre). For the Stokes problem, in addition to the loss behaviour for the original mesh, we also show the behaviour for a finer mesh.
Next, let us consider training errors, i.e. the error on the seen training data, as well as generalization errors, i.e. the errors seen on interpolating input data as well as on extrapolating input data. To evaluate the interpolation and extrapolation qualities, we test with unseen angles of attack within the training range, $\lambda\in\{5^{\circ},16.5^{\circ},30^{\circ},40^{\circ}\}$, and outside the training range, $\lambda\in\{47.5^{\circ},50^{\circ},55^{\circ}\}$, respectively. Figure 6 shows the average errors on the training data (”training”) as well as the generalization errors (”interpolation” and ”extrapolation”) in $L^{2}$- and $L^{\infty}$-norm for the different sizes of the training set. For clarity, we only report errors of the velocity in $x$-direction as the errors of velocity in $y$-direction and pressure behave similarly. It is clearly visible that preconditioning does not only heavily improve quality of the surrogate model on the training set itself, but also its inter- and extrapolation capabilities; in fact, the errors obtained without preconditioning are far beyond an acceptable range. The increase of the training errors for a growing size of the training set for the preconditioned problems is likely due to the fact that the same network architecture (i.e. the same number of degrees of freedom) is used to solve a increasingly complex problem. As expected, the generalization errors decrease with increasing training data.
@@
Figure 6: Averaged relative $L^{2}$- and $L^{\infty}$-errors of the variable $u_{x}$ predicted from FEM-based NNs trained with and without preconditioning for Stokes and Navier–Stokes ($\eta=1$) with several numbers of angles of attack #$\#\lambda_{\text{train}}$ and tested with angles of attack inside (interpolation) and outside (extrapolation) the training range. The dots represent the mean errors, the shadings correspond to the range in which the errors occurred.
@@ 4 Application: solving an inverse problem related to stationary 2D Navier–Stokes equations
In the following, we want to test this method. The details of the experiment are as follows: Figure 7 (right) shows the forward problem of predicting the pressure at the sensor locations for 17 angles of attack $\lambda$ that are equidistantly distributed in the range $[1^{\circ},45^{\circ}]$. Using this FEM-based NN as surrogate model, we generate 10 artificial and noisy pressure “measurements” by adding Gaussian noise with a standard deviation of $0.5$ to the respective pressure predictions evaluated at a reference angle of attack of $\lambda=5^{\circ}$. The “measurements” are shown in Figure 8.
@@
Figure 7: Left: Location of sensors at airfoil. Right: Pressure $p$ at sensor locations for different angles of attack $\lambda$ predicted by FEM-based NN.
@@
Figure 8: Left: Observed pressure measurements $y_{i}$ of two sensors. Right: Distribution of the angle of attack $\lambda$ of the airfoil generated with the MCMC algorithm.
@@ Appendix A Mesh-independent conditioning of the FEM layer for the preconditioned Stokes case
Let $\mathbf{Y}=\mathbf{Z}^{2}$ with $\mathbf{Z}$ as given in (21), then $\mathbf{Y}$ has the following annihilating polynomial

### v2

@@ 2.4 Example 2: Stationary Navier–Stokes equations in 2D
i.e., compared to the Stokes equations (7)-(8) in particular the non-linear convection term $u(\lambda)\cdot\nabla u(\lambda)$ has been added. Moreover, we have introduced the so-called kinematic viscosity $\eta>0$ as additional parameter. Discretization leads to
@@
with $b$ as before, $a$ as before but multiplied by $\eta$, and the trilinear form $c(z_{h};u_{h},v_{h})=\int_{\Omega}(z_{h}\cdot\nabla u_{h})\cdot v_{h}$ originating from the nonlinear term. We again use Taylor–Hood elements for the discretization in the non-linear case, as these are already necessary for a stable solution in the linear case. To formulate the corresponding nonlinear algebraic equation, we utilize Lagrangian basis functions as in (10) and obtain the discrete residual
@@
i.e., $\mathbf{C}(\mathbf{u})$ can be calculated by a contraction of the tensor $\mathbf{u}=\mathbf{L}^{-T}\tilde{\bf u}$ (or $\mathbf{u}=\tilde{\bf u}$ in the case of no preconditioning) with a 3D-tensor $\mathbf{\tilde{C}}$ that is independent of $\mathbf{u}$ and thus can be precomputed and reused. Since $\mathbf{\tilde{C}}$ is sparse, this can be done in a memory-efficient manner.
@@
$\mathbf{\hat{u}}_{i}:=\left\{\begin{array}[]{ll}\mathbf{u}_{i}&\text{if}\;x_{i}\notin\partial\Omega^{+}\cup\partial\Omega^{A},\\ u_{\textnormal{BC}}(x_{i})&\text{if}\;x_{i}\in\partial\Omega^{+}\cup\partial\Omega^{A},\end{array}\right.$ (19)
@@ 3 Numerical results
The respective neural networks and an FEM module as described in Sects. 2, 2.3, and 2.4 have been implemented in PyTorch [47]. Unless stated otherwise, the neural networks are plain vanilla feed forward networks consisting of five hidden layers with 50 neurons each and an output layer whose size is equal to the number of degrees of freedom of the discretization. When testing various hyperparameters, we found that small changes in the number of layers and number of neurons per layer did not have a major effect on the results. However, it turned out that a non-linear activation function which is not bounded is of great advantage to avoid local minima during optimization, which is why we use SELU [48] as the activation function except for the last, linear layer. To facilitate efficient training, we set the network weights with Xavier’s normal initialization [49] scaled by $3/4$. For the training, we use PyTorch’s implementation [^3] https://pytorch.org/docs/stable/generated/torch.optim.LBFGS.html [Accessed August 21, 2024] of L-BFGS. To make the results more comparable with “classical” implementations, we choose max_iter=1 to ensure that only one L-BFGS step is performed per iteration. Furthermore, we choose a line search within L-BFGS that seeks a solution that fulfills the strong Wolfe condition.
@@
[…]
Figure 4: Eigenvalues of the Hessian of the loss function, evaluated after training the FEM-based NNs with preconditioning (pre) and without preconditioning (nopre) for Stokes equations with an angle of attack $\lambda=1^{\circ}$. Eigenvalues are shown in gray, a kernel density estimate of the respective distribution functions in blue and orange. Note that the y-axis employs symmetric log axis scaling, i.e. logarithmic scaling with a linear scaling exception for values within $[-2,2]$.
@@ 3.1 The non-parametric case
Both the preconditioned and the non-preconditioned NNs are trained and evaluated for a single angle of attack $\lambda=1^{\circ}$ and 40000 L-BFGS iterations, for each: the Stokes problem and the Navier–Stokes equation. In the latter case, the solutions are computed with viscosities $\eta=1$ and $\eta=0.1$. Figure 3 shows that in all these cases preconditioning significantly speeds up the training in terms of a drastically reduced number of iterations. For the preconditioned cases we can even observe a stagnation of the residual at levels that are probably related to numerical accuracy; this is probably related to a similar reason as the effect described in the paragraph below (12) in Sect. 2.3. Figure 5 shows for the Navier–Stokes equation with $\eta=1.0$ that training is also much faster in terms of runtime because the very significant reduction in the number of iterations more than offsets the minimal increase in the runtime of a single iteration that is due to the few additional operations related to preconditioning.
@@
To further illustrate the effectiveness of the preconditioning, we computed the eigenvalues of the Hessian at the last iteration of the training of the Stokes equations, once with preconditioning and once without preconditioning. Note that computing the Hessian using the automatic differentiation machinery in PyTorch is possible, though possibly quite costly, and can only be employed for small problem sizes. The eigenvalues, and a kernel density estimate, for the two cases, with and without preconditioning, is plotted in Figure 4. It is clearly visible that even though our argumentation in Section 2 and A totally neglected the non-linearity introduced by the neural networks, the conclusion of clustered eigenvalues still holds. As noted earlier, we assume L-BFGS coupled with perfect line search to share the same characteristics as the CG method for a nearly quadratic optimization problem. Given that in our least-squares setting the cluster of zero eigenvalues can be neglected [52], the convergence of the CG algorithm for a quadratic optimization problem is determined by the number of remaining eigenvalue clusters [45]. Of course, this analysis is not exact, given that our problem is made non-linear by the neural network and potentially the advection term in the Navier–Stokes equations, and thus potentially deviates a lot from the aforementioned assumptions. Nonetheless, it clearly demonstrates the mechanisms at work, indicating the effectiveness of classical preconditioning techniques.
That the assumptions made in the design of the preconditioner are only approximately fulfilled, could not be more visible than in the high Reynolds number case, as illustrated in Figure 3. For $\eta=0.01$ the advection term clearly dominates, such that the chosen preconditioner, which takes only the diffusive part into account, performs poorly. The latter is clearly indicated by the increased number of iterations. Moreover, for higher Reynolds numbers techniques like, e.g, Streamline Upwind Petrov–Galerkin (SUPG) should be employed to further enhance the numerical properties [45, 53].
${10^{1}}$${10^{2}}$${10^{3}}$${10^{4}}$${10^{5}}$ime [s]${10^{-11}}$${10^{-8}}$${10^{-5}}$${10^{-2}}$${10^{1}}$${10^{4}}$ossNS, preNS, nopre
Figure 5: Loss values over time, in logarithmic scaling, when training FEM-based NNs with preconditioning (pre) and without preconditioning (nopre) for the Navier–Stokes (NS) equations with an angle of attack $\lambda=1^{\circ}$ and viscosity $\eta=1.0$. The initial construction times of the system matrices in (17) and (18) are shown in cyan, the training times of the iterations in purple.
@@ 3.2 The parametric case
Figure 7 shows the loss values during training for the different configurations under consideration. Even in this case, preconditioning is able to accelerate the decay of the loss by more than two orders of magnitude for both the Stokes and the Navier–Stokes equations. For the Stokes problem, we also performed the experiments on a finer mesh, which is obtained from the original one by a single uniform refinement and results in 28085 degrees of freedom. The behavior of the losses on the original and the finer mesh can be taken as an indication of mesh independence, which is consistent with our expectations from A and our discussion in the previous section.
@@
Figure 7: Loss values per L-BFGS iteration when training FEM-based NNs for multiple angles of attack #$\#\lambda_{\textnormal{train}}$ with preconditioning (pre) and without preconditioning (nopre). For the Stokes problem, in addition to the loss behavior for the original mesh, we also show the behavior for a finer mesh.
In the following, we will in particular consider the training errors, i.e., the error on the seen training data, as well as generalization errors: more specifically the errors seen on interpolating input data as well as on extrapolating input data. To evaluate the interpolation and extrapolation qualities, we test with unseen angles of attack within the training range, $\lambda\in\{5^{\circ},16.5^{\circ},30^{\circ},40^{\circ}\}$, and outside the training range, $\lambda\in\{47.5^{\circ},50^{\circ},55^{\circ}\}$, respectively. Figure 8 shows the average errors on the training data (”training”) as well as the generalization errors (”interpolation” and ”extrapolation”) in $L^{2}$- and $L^{\infty}$-norm for the different sizes of the training set. For clarity, we only report errors of the velocity in $x$-direction as the errors of velocity in $y$-direction and pressure behave similarly. It is clearly visible that preconditioning does not only heavily improve quality of the surrogate model on the training set itself, but also its inter- and extrapolation capabilities; in fact, the errors obtained without preconditioning are far beyond an acceptable range. The increase of the training errors for a growing size of the training set for the preconditioned problems is likely due to the fact that the same network architecture (i.e. the same number of degrees of freedom) is used to solve a increasingly complex problem. As expected, the generalization errors decrease with increasing training data.
@@
Figure 8: Averaged relative $L^{2}$- and $L^{\infty}$-errors of the variable $u_{x}$ predicted from FEM-based NNs trained with and without preconditioning for Stokes and Navier–Stokes ($\eta=1$) with several numbers of angles of attack #$\#\lambda_{\text{train}}$ and tested with angles of attack inside (interpolation) and outside (extrapolation) the training range. The dots represent the mean errors, the shadings correspond to the range in which the errors occurred for the given set of testing angles.
@@ 4 Application: Solving an inverse problem related to stationary 2D Navier–Stokes equations
In order to test this method, we choose the following setup: Figure 9 (right) shows the forward problem of predicting the pressure at the sensor locations for 17 angles of attack $\lambda$ that are equidistantly distributed in the range $[1^{\circ},45^{\circ}]$. Using this FEM-based NN as surrogate model, we generate 10 artificial and noisy pressure “measurements” by adding Gaussian noise with a standard deviation of $0.5$ to the respective pressure predictions evaluated at a reference angle of attack of $\lambda=5^{\circ}$. The “measurements” are shown in Figure 10.
@@
Figure 9: Left: Location of sensors on airfoil. Right: Pressure $p$ at sensor locations for different angles of attack $\lambda$ predicted by FEM-based NN.
@@
Figure 10: Left: Observed pressure measurements $y_{i}$ of two sensors. Right: Corresponding posterior distribution of the angle of attack $\lambda$ of the airfoil, as generated with the MCMC algorithm.
@@ Appendix A Mesh-independent conditioning of the FEM layer for the preconditioned Stokes case
Consider $\mathbf{Y}=\mathbf{Z}^{2}$ with $\mathbf{Z}$ as given in (21), then $\mathbf{Y}$ has the following annihilating polynomial
@@ Acknowledgements
Last but not least, the authors want to thank the anonymous reviewers for their valuable comments.